Grade 12th passPhysical Chemistry

Fe(OH)2 is a dibasic acid with Kb1=10^-4 and Kb2=2.5×10^-6.what is the concentration of Fe(OH)2 in 0.1M Fe(NO3)2 solution ? Ans:10^-10

Pawan joshi
7 Years agoGrade 12th pass

1 Answer

Arun
7 Years ago
kb1=(10)- 4 and kb2=2.5*(10)- 6 
Convergence of ferrous nitrate = 0.1 M 
At that point convergence of ferrous hydroxide = ? 
Kb = ([HB+] [OH¯])/[B] 
10 - 4 = 0.1 x 2.5* 10 ^ - 6 concn 
10-4=0.25 * 10^6 conc 
Concn = 0.25 * 10^6/10-4 
Concn = 0.0005mole/l 
 
Hope it helps